### v1

# CT-Mamba: A Hybrid Convolutional State Space Model for Low-Dose CT Denoising

Abstract
Low-dose CT (LDCT) significantly reduces the radiation dose received by patients, thereby decreasing potential health risks. However, dose reduction introduces additional noise and artifacts, adversely affecting image quality and clinical diagnosis. Currently, denoising methods based on convolutional neural networks (CNNs) face limitations in long-range modeling capabilities, while Transformer-based denoising methods, although capable of powerful long-range modeling, suffer from high computational complexity. Furthermore, the denoised images predicted by deep learning-based techniques inevitably exhibit differences in noise distribution compared to Normal-dose CT (NDCT) images, which can also impact the final image quality and diagnostic outcomes. In recent years, the feasibility of applying deep learning methods to low-dose CT imaging has been demonstrated, leading to significant achievements. This paper proposes CT-Mamba, a hybrid convolutional State Space Model for LDCT image denoising. The model combines the local feature extraction advantages of CNNs with Mamba’s global modeling capability, enabling it to capture both local details and global context. Additionally, a Mamba-driven deep noise power spectrum (NPS) loss function was designed to guide model training, ensuring that the noise texture of the denoised LDCT images closely resembles that of NDCT images, thereby enhancing overall image quality and diagnostic value. Experimental results have demonstrated that CT-Mamba performs excellently in reducing noise in LDCT images, enhancing detail preservation, and optimizing noise texture distribution, while demonstrating statistically similar radiomics features to those of NDCT images (p $>$ 0.05). The proposed CT-Mamba demonstrates outstanding performance in LDCT denoising and holds promise as a representative approach for applying the Mamba framework to LDCT denoising tasks.
Index Terms:
Low-Dose CT, Denoising, State Space Model, Mamba, Noise Power Spectrum, Radiomics
I Introduction
Computed tomography (CT) is an essential imaging technique in clinical practice, providing crucial anatomical information that aids physicians in making appropriate medical decisions. However, frequent CT scans can significantly increase the radiation dose received by patients, particularly in radiation therapy, where multiple scans are often required to achieve precise tumor localization and minimize damage to organs at risk. Excessive radiation may harm the patient, impact organ function, increase the incidence of radiation-related diseases linked to genetic damage, and ultimately reduce the patient’s quality of life. Therefore, reducing the radiation dose of CT imaging has become a focus of attention. Low-dose CT (LDCT) is an effective method for reducing patient radiation exposure [1]. In clinical practice, to acquire LDCT images, it is common to reduce the tube current to decrease the flux of the electron beam that generates X-ray photons, and/or reduce the number of projections during CT procedures. While these methods reduce radiation dose, they also introduce additional noise and artifacts, which decrease the overall quality of the image. If these issues are not effectively addressed, they will significantly impact the application of LDCT in various clinical scenarios. How to obtain high-quality CT images that are comparable to normal-dose CT (NDCT) images and meet the clinical utilities from low-dose scanning protocols has become a long-standing and practically significant problem in the field of CT imaging.
To address this issue, several LDCT imaging algorithms have been proposed, which can be roughly divided into the following three categories [2, 3]: sinogram-based preprocessing, iterative reconstruction, and image post-processing. Sinogram-based preprocessing methods directly target the raw data acquisition stage and designs specific filters for processing CT projection data obtained through low-dose X-rays. Typical methods include bilateral filtering [4], structure-adaptive filtering [5], and penalized weighted least squares [6]. These denoising methods combine physical characteristics with photon statistical characteristics, making them relatively simple and easy to implement. However, these methods rely heavily on high-quality original projection data, limiting their ability to effectively restore undersampled or missing signals, and their practical application is limited by the difficulty in obtaining complete sinogram data. Iterative reconstruction methods [7] typically reconstruct images by iteratively optimizing an objective function. It usually alternates between forward and backward projections in the projection domain and the image domain until the objective function is minimized based on the convergence criterion. However, this method typically requires access to the raw data and bears a high computational cost. In addition, the quality of image reconstruction heavily relies on the precise setting of the objective function and hyperparameters. These limitations hinder the application of iterative reconstruction methods in clinical practice.
In contrast, image post-processing methods primarily address noise and artifacts within the image domain. Typical methods include non-local means filtering algorithm [8], and block matching algorithm [9]. Compared to the first two types of methods, post-processing has the advantage of not requiring original projection data from the CT vendor, making it easier to integrate into clinical CT imaging workflows.
In recent years, with the continuous improvement of computer performance, and the rapid development of cloud computing and distributed computing, deep learning-based image post-processing methods have achieved remarkable achievements in the field of LDCT denoising. Deep learning can automatically learn complex patterns, offering distinct advantages over traditional image post-processing methods. Currently, two popular network architectures, convolutional neural networks (CNNs) [10] and Transformers [11], are widely applied in LDCT denoising. Furthermore, hybrid models combining these two architectures, along with other neural network frameworks, are continually being developed.
CNN-based methods effectively extract image features through convolutional layers, and have achieved remarkable success in tasks such as image restoration [12], object detection [13], and image segmentation [14]. In the field of LDCT denoising, several representative CNN-based models have been proposed [15, 16, 17, 18, 19, 20, 21]. However, CNN frameworks are limited by their lack of long-range modeling capabilities, making it difficult to capture the global contextual information in images [22]. Transformers were originally designed for natural language processing, but have now been successfully adapted for image processing, such as Vision Transformer (ViT) [23] for image recognition and SwinTransformer [24] for a general backbone in various vision tasks. By treating images as sequences of patches, Transformers has better capabilities to capture global information. However, they are computationally expensive, and this is because the self-attention mechanism scales quadratically with the input size [25]. This characteristic creates a serious bottleneck, particularly in high-resolution applications and edge computing scenarios. To overcome these limitations, researchers are working to improve these models.
In this context, the structured state space sequence models (S4) have attracted extensive attention due to their efficient performance in long-range modeling [26]. S4 optimizes its internal state representation and information transmission mechanism, significantly reducing computational and memory requirements, thus showing great potential in large-scale sequence modeling tasks. However, despite its breakthroughs in long-range modeling and computational efficiency, S4 still faces limitations in contextual reasoning and flexible information selection. To further enhance the model’s performance in complex scenarios, the Mamba model (S6) is developed [27]. Mamba integrates a selection mechanism into S4, allowing the model to selectively propagate or discard information along the sequence. This improvement boosts the model’s contextual reasoning abilities while maintaining the computational efficiency of S4.
Currently, Mamba has been rapidly adopted in computer vision tasks. It unfolds image patches into sequences along the horizontal and vertical dimensions of the image, and performs bi-directional scanning along these two directions. This bidirectional scanning method enables VMamba to effectively capture both global and local information within images. At present, some studies have employed the network architectures based on visual Mamba [28, 29, 30, 31, 32]. Additionally, applications in LDCT image denoising have also been reported [33]. The classic visual Mamba scanning methods, such as horizontal and vertical scanning, can cause certain adjacent pixels in the spatial domain to be spaced too far apart when unfolded into a sequence. This separation can undermine Mamba’s performance in visual tasks, particularly in medical imaging, where capturing fine structures and lesions is critical. To address this issue, we have improved the scanning method to ensure the spatial correlation between pixels in the image. Furthermore, by combining Mamba with CNN, we harness CNN’s strengths in local feature extraction along with Mamba’s global modeling capability, allowing the model to excel in handling the complex task of LDCT image denoising.
Current researches also ignore an important issue: the inability to quantitatively characterize noise distribution in predicted images. This may lead to problems such as excessive smoothing, misjudgment of lesions and key structures, and loss of spatial resolution. The Noise Power Spectrum (NPS) can quantitatively describe the noise texture in an image [34]. Therefore, this paper also utilizes NPS to guide deep learning LDCT denoising tasks, ensuring that the predicted images more accurately restore the noise distribution present in NDCT images, thereby improving overall image quality and diagnostic value.
The contributions of this study are summarized as follows:
1. We propose CT-Mamba, a hybrid convolutional state space model designed for LDCT image denoising. This model integrates the multi-scale analysis capability of wavelet transform, the powerful local feature extraction capacity of CNN, and the long-range dependency modeling strengths of Mamba, enabling comprehensive feature capture within the images.
2. We propose the Coherence Z-Scan State Space Block (CZSS), which introduces an innovative spatially coherent “Z”-shaped scanning scheme. This approach ensures spatial continuity between adjacent pixels in the image, enhancing the model’s ability to preserve details and improve denoising effectiveness.
3. To ensure that the denoised LDCT images can closely restore the noise texture of NDCT, this study also designs a Mamba-driven Deep NPS Loss (Deep NPS Loss).
4. Using radiomics, the proposed CT-Mamba was evaluated by comparing the statistical distribution of radiomics features in different organs between denoised LDCT images and NDCT images, as well as the mean absolute error of pairwise features.
Figure 1: (a) The overall architecture of the proposed CT-Mamba. (b) The structure of the Multi-Scale Coherence Mamba architecture (MSC-Mamba) in CT-Mamba.
II Medthod
II-A Preliminary: State Space Model
The state space model (SSM) is a mathematical description of a dynamic system which maps a one-dimensional input function or sequence $x(t)\in{{\mathbb{R}}^{L}}$ to an output $y(t)\in{{\mathbb{R}}^{L}}$ through hidden states $h(t)\in{{\mathbb{R}}^{N}}$. From a mathematical perspective, this process can be represented by a linear Ordinary Differential Equation (ODE):
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t)$ (1)
$y(t)$ $=\mathbf{C}h(t),$
where $\mathbf{A}\in{{\mathbb{R}}^{N\times N}}$ represents the state matrix, while $\mathbf{B}\in{{\mathbb{R}}^{N\times 1}}$ and $\mathbf{C}\in{{\mathbb{R}}^{1\times N}}$ denote the projection parameters.
In order to integrate SSM into deep learning, the linear ordinary differential equation mentioned above needs to be discretized. Typically, the discretization of the SSM uses the Zero-Order Hold (ZOH) method. By incorporating a time scale parameter $\Delta$, the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ are transformed into discrete parameters $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$, which can be defined as follows:
$\overline{\mathbf{A}}=\exp(\Delta\mathbf{A})$ (2)
$\overline{\mathbf{B}}={{(\Delta\mathbf{A})}^{-1}}(\exp(\Delta\mathbf{A})-I)\cdot\Delta\mathbf{B}$
After discretization, Equation (1) can be reformulated as:
${h}^{\prime}(t)=\overline{\mathbf{A}}h(t)+\overline{\mathbf{B}}x(t)$ (3)
$y(t)=\mathbf{C}h(t)$
Equation (3) results in a discretized state space model that can be computed recursively. However, due to its sequential nature, this discretized recursive SSM is impractical for training. To enable efficient parallelized training, this recursive process can be reformulated as a convolution for computation:
$\overline{\mathbf{K}}=(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{A}}\overline{\mathbf{B}},\ldots,\mathbf{C}{{\overline{\mathbf{A}}}^{L-1}}\overline{\mathbf{B}})$ (4)
$y=x*\overline{\mathbf{K}},$
where $\overline{\mathbf{K}}\in{{\mathbb{R}}^{L}}$ represents a structured convolutional kernel, $L$ denotes the length of the input sequence $x$, and * represents the convolution operation.
The recently proposed Mamba model makes further improvements by introducing a selective scanning mechanism. This allows parameters $\mathbf{B}$, $\mathbf{C}$ and $\Delta$ to be dynamically adjusted based on the input $x$ and contextual information. As a result, Mamba can model complex temporal dynamics more effectively, as the model can adapt to the evolving characteristics of the input data.
II-B Framework Overview
In this paper, we propose CT-Mamba, a hybrid convolutional state space model designed for LDCT image denoising. This model integrates the multi-scale analysis capability of wavelet transform, the powerful local feature extraction ability of CNN, and Mamba’s long-range dependency modeling advantages, enabling it to comprehensively capture image features, as shown in Fig. 1(a).
In the initial stage of CT-Mamba, we designed a lightweight Progressive Feature Extraction Network (PFEN). This network extracts features at two levels, gradually capturing primary spatial features of LDCT images while incorporating raw information from LDCT images to provide richer spatial information for subsequent wavelet domain processing. Then, we applied a first-level wavelet transform to decompose the spatial information into a low-frequency component, a horizontal high-frequency component, a vertical high-frequency component, and a diagonal high-frequency component. This helps CT-Mamba better capture different frequency features in LDCT images. At the end of the network, we also designed a lightweight Progressive Feature Fusion Network (PFFN), which performs feature recombination across two branches. This enables efficient integration and refinement of features extracted from the wavelet domain. These two spatial networks work in tandem, effectively capturing and processing spatial information while significantly reducing the model’s computational complexity. Through combined processing in both the spatial and wavelet domains, the quality of the output image is notably enhanced.
After the wavelet transform, although the components in the horizontal, vertical, and diagonal directions are decomposed independently, they still exhibit certain correlations. For example, the edges or texture details of an object may simultaneously manifest in the horizontal, vertical, and diagonal directions. We believe that the correlation between high-frequency information in the diagonal direction and that in the horizontal or vertical directions is generally stronger than the correlation between horizontal and vertical information. Based on this insight, we designed a feature fusion strategy incorporating a High-Frequency Feature Extraction Network (HFEN), as shown in Fig. 2. The HFEN is equipped with multi-scale feature extraction, cross-scale feature fusion, and channel compression capabilities, which significantly enhances its ability to extract features from frequency information in different directions. By fusing high-frequency information from the horizontal and vertical directions into the diagonal direction, and vice versa, this strategy facilitates effective interactions among high-frequency components. This enables the network to better capture complex structural features within the image, thereby improving overall performance.
Figure 2: The structure of High-Frequency Feature Extraction Network (HFEN).
We designed a Multi-Scale Coherence Mamba architecture (MSC-Mamba) to learn the features of each frequency branch obtained by wavelet transform decomposition, as shown in Fig. 1(b). MSC-Mamba captures and fuses frequency features learned from three different scales (32×32, 16×16, and 8×8) by integrating Coherence Z-Scan State Space Block (CZSS). At lower scales, MSC-Mamba effectively captures global feature information, while at higher scales, it focuses more on local details. By combining feature learning across different scales, MSC-Mamba effectively ensures spatial coherence, providing superior denoising performance. Additionally, inspired by [35], we designed a Frequency Domain Attention Module (FDAM) based on Fourier transform to enhance the low-frequency features in the wavelet domain, as shown in Fig. 3. Each grid in the extracted Fourier spectrum contains global information on the low-frequency features, enabling efficient modeling of long-range spatial dependencies in the frequency domain. Enhanced by FDAM, richer low-frequency characteristics are provided for MSC-Mamba located in the low-frequency branch, thereby optimizing the modeling of overall structural information and improving CT-Mamba’s performance.
Figure 3: The structure of Frequency Domain Attention Module (FDAM).
II-C Coherence Z-Scan State Space Block
The Coherence Z-Scan State Space Block (CZSS) is a novel feature extraction unit specifically designed for tasks like medical imaging. Classic visual Mamba models, such as VMamba, which excel in natural image processing, utilize a row-column scanning order with bi-directional scanning in both horizontal and vertical directions. However, such scanning methods can overly separate some adjacent pixels when unfolding an image into a sequence. For instance, in horizontal scanning, there is a substantial distance between the end of one row and the beginning of the next, weakening spatial connections and subsequently hindering the capture of fine structures. To address this issue, inspired by the Zigzag scanning method in JPEG image compression theory, Z-SSM adopts a similar Z-shaped scanning approach to maintain spatial continuity between each adjacent pixel in the image, as shown in Fig. 4. Z-SSM employs multi-directional scanning and combined with a selective state space model SSM (S6) for modeling, ultimately merging the sequence and restoring it to a two-dimensional structure. This design ensures that subtle lesions and detailed information within the image are not overlooked, thereby enhancing the denoising effect of LDCT images.
Figure 4: Illustration of the structure of Z-SSM. The Z-shaped scanning method is adopted to ensure spatial continuity between each adjacent pixel in the image.
A residual block with two convolutional layers is added to the front of CZSS can locally maintain the frequency-spatial correlation of input features and efficiently transfer local contextual information, thereby enabling more effective utilization and further optimization of these features, as represented by “Res” in Fig. 1(b). The CZSS begins by standardizing the input data through layer normalization, improving training stability. Next, a linear transformation adjusts the data dimensions, followed by depthwise separable convolution, which operates on each input channel independently, reducing parameter count and focusing on feature extraction. After convolution, the SiLU activation function is applied, and its smooth gradient properties help the model to stably learn nonlinear features, thereby enhancing its ability to capture complex patterns. Then, Z-SSM performs Z-shaped scanning across four paths, after which the outputs are merged and restructured into a 2D form. Layer normalization is then applied, and the SiLU activation is used to adjust deep features. Finally, a linear layer is applied with DropPath to enhance robustness and mitigate overfitting. CZSS finds a good balance between computational efficiency and feature extraction capability, allowing for stacking more CZSS blocks under similar depth constraints.
II-D Deep Noise Power Spectrum Loss
Noise texture is highly sensitive to changes in subjective visual quality. NPS, based on Fourier transform, can characterize the correlation between pixels from a frequency domain perspective, thereby providing a quantitative description of image noise texture. In CT imaging, noise originates from various factors, including but not limited to the signal acquisition system, radiation dose, and computational algorithms.
In LDCT denoising tasks, neglecting the optimization of noise texture can severely impact the quality of denoised images. Therefore, quantitative analysis of the NPS is crucial for optimizing and evaluating image quality. In this study, we utilize NPS to guide the denoising of LDCT images for the first time, aiming to more accurately restore the noise texture in the denoised images. The NPS in this study is defined as follows:
$\text{NP}{{\text{S}}_{2D}}(u,v)=\frac{{{p}_{x}}{{p}_{y}}}{{{N}_{x}}{{N}_{y}}}{{\left|\text{DFT}(\text{noise(}m,n))\right|}^{2}},$ (5)
[…]
Since it is difficult for medical images to separate noise from a single image, we designed a dual-branch structure optimized based on NPS, as shown in Fig. 5(a). By subtracting the NDCT image from the LDCT image, we obtain reference noise (denoted as $\text{nois}{{\text{e}}_{\text{GT}}}$, $\text{nois}{{\text{e}}_{\text{GT}}}={{I}_{\text{LDCT}}}-{{I}_{\text{NDCT}}}$), and by subtracting the model-predicted image from the LDCT image, we obtain the predicted noise (denoted as $\text{nois}{{\text{e}}_{\text{Pred}}}$, $\text{nois}{{\text{e}}_{\text{Pred}}}={{I}_{\text{LDCT}}}-{{I}_{\text{Pred}}}$). As each image can be abstracted as a combination of signal and noise, so the difference between these two noise components lies solely in the noise present in the model-predicted image versus the NDCT image. By analyzing NPS based on the pair of noise, we can guide the model to generate images with a noise distribution closer to that of NDCT images.
In the dual-branch structure, we designed an enhanced U-shaped network (u-Feature Net) to extract features from $\text{nois}{{\text{e}}_{\text{GT}}}$ and $\text{nois}{{\text{e}}_{\text{Pred}}}$, as shown in Fig. 5(b). Low-frequency information is crucial in noise modeling, particularly for maintaining the stability of the overall noise structure and texture. Effective capture of low-frequency features is thus essential for improving noise texture. To address this, we designed a Mid and Low Frequency Feature Enhancement Connection (MLF-FEC) within u-Feature Net, specifically reinforcing the network’s processing of mid- and low-frequency information. Building on this, we applied an NPS transformation to the extracted noise features and incorporated the long-range modeling capability of CZSS to accurately capture the global noise feature distribution across the full frequency. This design effectively guides CT-Mamba in optimizing noise distribution.
After the full-band modeling of CZSS, the extracted $\text{nois}{{\text{e}}_{\text{GT}}}$ and $\text{nois}{{\text{e}}_{\text{Pred}}}$ features are converted into one-dimensional radial NPS (Radial NPS) to reduce complexity while retaining critical frequency information. The radialization principle is shown in Fig. 6. We then use L1 loss to measure the difference between the two radial NPS signals and design a correlation loss based on the Pearson correlation coefficient to guide model optimization. This approach ensures that the noise texture in the model’s denoised images more closely resembles that of NDCT images, thereby improving overall image quality and diagnostic value.
Figure 5: Detailed structure of the dual-branch Deep NPS Loss and its key components. (a) Structure of the Deep NPS Loss. (b) Detailed structure of the u-Feature Net.
Figure 6: Schematic diagram of the radial transformation process of a 2D NPS signal.
II-E Overall Objective
The loss function measures the difference between the model’s predicted output (the denoised LDCT image) and the ground truth (NDCT image). While the model architecture determines the model’s complexity, the loss function controls how denoising characteristics are learned from the training dataset. To recover high-quality denoised CT images from LDCT, we combine different loss components into a hybrid objective function:
1) Pixel-level L1 Loss: Compared with L2 loss (mean square error, MSE), L1 loss does not overly penalize large errors between the predicted and NDCT images, so it can alleviate the blurriness and unnatural effects caused by L2 loss to a certain extent. In our LDCT denoising task, it is defined as (6):
[…]
2) Deep NPS Loss: This loss component ensures that the noise texture in the denoised image predicted by CT-Mamba is closer to that of the NDCT image. It is represented in conjunction with the description in Section 2.4 as (7):
${{\mathcal{L}}_{\text{deep-NPS}}}(x,y,\hat{y})=$ $\,{{\gamma}_{1}}\cdot{{\left\|\text{NP}{{\text{S}}_{\text{radial}}}(\phi(x,y)-\phi(x,\hat{y}))\right\|}_{1}}$ (7)
$+{{\gamma}_{2}}\cdot\left(1-\rho\left(\text{NP}{{\text{S}}_{\text{radial}}}(\phi(x,y)),\text{NP}{{\text{S}}_{\text{radial}}}(\phi(x,\hat{y}))\right)\right),$
[…]
3) Feature-Level Perceptual Loss: Perceptual loss captures information at the feature level using a pre-trained deep network. This loss is based on ResNet50 [36], pre-trained on the ImageNet dataset, with its weights frozen during training. Since the pre-trained ResNet50 accepts color images as input, we replicated the grayscale CT images across three channels to fit this input requirement. we extracted feature maps after its four main stages and used L2 loss to measure the similarity between the CT-Mamba output and the ground truth image. The perceptual loss is defined as (8):
$\mathcal{L}_{\text{perceptual}}(y,\hat{y})=\sum_{i=1}^{4}\|f_{i}(y)-f_{i}(\hat{y})\|_{2},$ (8)
[…]
In summary, the overall objective function is the weighted sum of each loss term (9):
${{\mathcal{L}}_{\text{total}}}={{\lambda}_{1}}\cdot{{\mathcal{L}}_{\text{L1}}}+{{\lambda}_{2}}\cdot{{\mathcal{L}}_{\text{deep-NPS}}}+{{\lambda}_{3}}\cdot{{\mathcal{L}}_{\text{perceptual}}},$ (9)
where ${{\lambda}_{2}}=({{\gamma}_{1}},{{\gamma}_{2}})$.
III EXPERIMENT DESIGNS AND RESULTS
III-A Datasets
In this work, we used a publicly released patient dataset for the 2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge. This dataset contains paired low-dose and normal-dose abdominal images with a slice thickness of 1 mm for 10 patients. The quarter-dose LDCT images were generated by adding Poisson noise to the projection data of the NDCT images to mimic a noise level that corresponded to 25% of the full dose. We used data from 9 patients for model training, totaling 5,410 image pairs, while data from the remaining patient (L506) was used for testing model performance, totaling 526 pairs. Additionally, using the method in [37], we simulated paired low-dose and normal-dose images for 20 patients from the AMOS dataset (referred to as the Simulator dataset) for model training and validation. The low-dose images also correspond to a quarter of the normal dose. We used data from 19 patients for training, totaling 1,872 image pairs, and data from 1 patient (AMOS-0033) for testing model performance, totaling 106 pairs.
III-B Implementation details
The proposed CT-Mamba model is implemented in PyTorch and trained in an environment equipped with an NVIDIA RTX 3090 Ti 24G GPU. The optimizer used is AdamW, with parameters set to ($\beta 1$, $\beta 2$) = (0.9, 0.999) and a weight decay of 0.02. The initial learning rate is set to 1e-3 and is gradually reduced to 1e-6 using cosine annealing. During training, each image is randomly cropped into four 64×64 patches as input for each batch, with the batch size of 8. To balance the numerical relationship between the deep NPS-loss and the L1 loss and ensure training stability, the deep NPS-loss is not introduced during the first 10 epochs. Afterward, the hyperparameters of corresponding weights for the deep NPS loss are set to ${{\lambda}_{2}}$ = (1e-4, 1e-2). The perceptual loss weight ${{\lambda}_{3}}$ is set to 1e-2. The training process lasts for 250 epochs, with the final model selected based on the optimal loss achieved.
Figure 7: A set of slice prediction results and difference images obtained from the Mayo dataset L506 patient. The enlarged ROI (blood vessel) marked by the red dashed rectangle is located in the lower right corner of the image.
Figure 8: A set of slice prediction results and difference images obtained from the Mayo dataset L506 patient. The enlarged ROI (low-attenuation lesion, tissue edge) marked by the red dashed rectangle is located in the lower right corner of the image.
III-C Experiment result and design
To demonstrate the effectiveness of the proposed CT-Mamba, we selected several advanced representative methods for comparative experiments, including EDCNN [38], RED-CNN [39], Uformer [40], CTformer [41], and VM-UNet [42]. EDCNN and RED-CNN are CNN-based methods, Uformer and CTformer are Transformer-based methods, and VM-UNet is based on the Mamba framework.
III-C1 Visual evaluation
This section presents the visual validation results of the CT-Mamba model compared to various baseline methods. In the Mayo dataset, we selected two representative slices, each containing the predicted results from each method along with their difference images relative to the NDCT images, to validate the comprehensive performance of our proposed method in LDCT processing, as shown in Fig. 7 and Fig. 8. To further assess detail representation, an enlarged view of the region of interest (ROI) which is marked by a red dashed rectangle is provided in the lower right corner of each image, along with the PSNR/SSIM quantitative results for each method displayed in the lower left corner.
From the results shown in Fig. 7 and Fig. 8, all methods exhibit a certain degree of noise and artifact removal capability. However, compared to other methods, our proposed approach achieves the optimal processing performance, demonstrating excellent denoising competence in terms of detail preservation and artifact suppression. In the ROIs, we observe that CT-Mamba enhances the representation of fine structures (such as blood vessels), as shown in Fig. 7; low-attenuation lesions are more clearly visible while maintaining well-defined tissue edges, as shown in Fig. 8.
In Fig. 9, we also present the denoising results and difference images of a representative slice from the Simulator dataset, further validating the outstanding denoising performance and generalization capability of CT-Mamba across different datasets. The upper right corner of each predicted image shows the corresponding difference image.
Figure 9: A set of slice prediction results and difference images obtained from the Simulator dataset AMOS-0033 patient, with the difference images located in the upper right corner.
We set the window level of the predicted results to 40 HU and the window width to 400 HU (i.e., a range from -160 HU to 240 HU). For the difference images, the window level is set to 0 HU with a window width of 200 HU (i.e., a range from -100 HU to 100 HU).
III-C2 Quantitative analysis
Our method not only demonstrates significant optimization effects in subjective visual evaluation but also achieves excellent performance across multiple quantitative metrics. Most studies use common quantitative metrics in image processing to report their results, such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure(SSIM), and Root Mean Square Error (RMSE). However, existing studies have indicated that these metrics do not sufficiently reflect clinical relevance [43]. Although human readers are considered as the gold standard for evaluating medical images, conducting multi-reader studies is time-consuming and costly. The research by Eulig et al. indicates that, compared to quantitative metrics like SSIM and PSNR for CT and MR images, various metrics including Visual Information Fidelity (VIF) have a higher correlation with human ratings [44]. Therefore, this study employs PSNR, SSIM, RMSE, and VIF for comprehensive evaluation. Additionally, in next section, we conducted a radiomics analysis to extract and compare the similarity of radiomic features.
In Table. I and Table. II, we present the average quantitative results for PSNR, SSIM, RMSE, and VIF for patient L506 from the Mayo dataset and patient AMOS-0033 from the Simulator dataset, respectively. A comprehensive evaluation of the PSNR, SSIM, RMSE, and VIF metrics shows that CT-Mamba model outperforms the comparison methods across all quantitative indicators.
TABLE I: The average quantitative results of different methods for patient(L506) in the Mayo test dataset.
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ VIF $\uparrow$
LDCT 42.3509 0.9471 31.9809 0.2965
EDCNN 46.6959 0.9819 19.1911 0.3578
REDCNN 46.1284 0.9799 20.5049 0.3406
Uformer 44.4342 0.9655 24.8988 0.3192
CTformer 46.2254 0.9770 20.3881 0.3471
VM-Unet 46.5672 0.9824 19.4795 0.3606
CT-Mamba 47.2825 0.9842 17.9850 0.3735
TABLE II: The average quantitative results of different methods for patient(AMOS-0033) in the Simulator test dataset.
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ VIF $\uparrow$
LDCT 38.7673 0.9099 35.5206 0.3573
EDCNN 50.3082 0.9931 9.2005 0.5186
REDCNN 49.7609 0.9921 9.8209 0.5098
Uformer 44.9416 0.9795 17.1556 0.4241
CTformer 43.1905 0.9673 21.4047 0.4063
VM-Unet 48.0132 0.9917 11.9833 0.4880
CT-Mamba 51.2054 0.9942 8.2936 0.5444
TABLE III: The p-values calculated using the Wilcoxon rank-sum test for each model. AP-value of under 0.05 indicates statistical difference. Pairs with statistical differences are marked by *.
Method Aorta Right Kidney Liver Stomach Small Bowel Urinary Bladder
LDCT 0.2774 0.0787 0.1626 0.7841 0.7116 0.6320
EDCNN 0.0557 0.0557 0.3793 0.4701 0.7519 0.5830
REDCNN 0.0979 0.4080 0.7959 0.9191 0.9893 0.6018
Uformer 0.0979 0.1208 0.0386* 0.2137 0.4772 0.1766
CTformer 0.1960 0.1626 0.3544 0.6198 0.6498 0.3040
VM-Unet 0.0016* 0.3519 0.8361 0.9801 0.8286 0.6741
CT-Mamba 0.5014 0.7960 0.6872 0.8948 0.8857 0.8032
TABLE IV: Number of similarity ratio which are higher for each model compared to the LDCT volumes out of the 93 features.
Method Aorta Right Kidney Liver Stomach Small Bowel Urinary Bladder
EDCNN 6 1 6 32 61 8
REDCNN 5 1 5 33 81 12
Uformer 10 5 5 62 76 10
CTformer 8 1 8 36 63 7
VM-Unet 9 5 6 23 67 13
CT-Mamba 4 2 7 11 4 10
III-C3 Radiomics reserach
Radiomics holds the potential to transform digital medical images into quantitative features that reveal underlying pathology, with promising applications in tumor classification and patient outcome prediction [45].In this study, the TotalSegmentator [46] extension in 3D Slicer with a pretrained nnU-Net [47] was utilized for automatic segmentation of target organs (aorta, right kidney, liver, stomach, small bowel, and urinary bladder), followed by manual adjustments to ensure accuracy. For patient L506 from the Mayo dataset, 93 first-order radiomics features were extracted for each target organ from the volumes using Pyradiomics (ver. 3.0). An ideal LDCT denoising algorithm is expected to produce LDCT volumes with statistically similar radiomics features to those of NDCT volumes. The following two tests were designed to show statistical similarity:
(1) Statistical Distribution of First-Order Features: In statistical analysis, the Wilcoxon rank-sum test is commonly used to assess whether two datasets share the same distribution. For each volume, the Wilcoxon rank-sum test was employed to evaluate the consistency of radiomics features distribution between the denoised images and the reference NDCT (p $>$ 0.05). In Table III, we show the p-values for each model across the target organs (aorta, right kidney, liver, stomach, small bowel, and urinary bladder). A higher p-value indicates greater similarity to the radiomics features distribution of the NDCT. As shown in Table III, while VM-UNet performed well for the liver and stomach, its performance for the aorta was poor, exhibiting significant differences in radiomics features distributios compared to the NDCT. In contrast, the proposed CT-Mamba model demonstrated superior performance across multiple organs, including the aorta, right kidney, and urinary bladder. Furthermore, its radiomics features distribution closely aligned with those of the NDCT for all other target organs. Notably, for the aorta, the CT-Mamba outperformed all other models.
(2) Pairwise Feature Similarity Ratio: To quantitatively assess the similarity between the radiomics features of the target organs for all models and those of the NDCT, pairwise comparisons were performed for the 93 first-order features. The mean absolute error (MAE) between the features of each model and the corresponding NDCT features (10):
$R_{i}=\left|\frac{V_{\text{NDCT}}-V_{i}}{V_{\text{NDCT}}}\right|,$ (10)
where $i$ was the different models, and $V$ was the values of the selected radiomics features. A lower similarity ratio $R_{i}$ indicates higher similarity between the features of the model and those of the NDCT. For each model, the $R$ across the target organs were compared against $R_{\text{LDCT}}$, and the number of features with $R$ exceeding $R_{\text{LDCT}}$ was recorded.
As shown in Table IV, the proposed CT-Mamba demonstrated superior performance across all target organs. Notably, for the small bowel, where other models performed poorly, CT-Mamba achieved the best performance, further highlighting its robustness and effectiveness. The radiomics analysis indicate that the proposed CT-Mamba effectively preserves the shape and texture features of multiple organs within the LDCT images, demonstrating its potential value for clinical applications.
III-C4 Ablation study
To evaluate the effectiveness of the key components of CT-Mamba, we designed three ablation experiments and conducted quantitative validation on patient L506 from the Mayo dataset:
(A1) Effectiveness of deep NPS loss: Excluding deep NPS-loss during training.
(A2) Effectiveness of the Frequency Domain Attention Module: Excluding the FDAM module.
(A3) Effectiveness of the Scanning Method: Replacing the Coherence ”Z” Scan with the classic Visual Mamba scanning method.
(M) Our complete model.
The quantitative results of each ablation experiment are summarized in Table. V. The results indicate that the complete CT-Mamba model achieves the best performance across all metrics, further validating the significant contribution of each component to the overall performance of model.
TABLE V: Quantitative comparison of different ablation experiments for patient(L506) in the Mayo test dataset.
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ VIF $\uparrow$
LDCT 42.3509 0.9471 31.9809 0.2965
A1 47.0101 0.9830 18.5470 0.3679
A2 47.2301 0.9841 18.0940 0.3723
A3 47.2245 0.9841 18.1040 0.3728
M 47.2825 0.9842 17.9850 0.3735
To further investigate the impact of deep NPS loss of CT-Mamba in terms of noise texture distribution, we conducted an independent NPS analysis of the slice shown in Fig. 8, focusing on the uniform ROI area marked in Fig. 10(a), with a size of 32×32. By comparing the radial NPS characteristics of the ROI area under conditions without deep NPS loss (A1) and with deep NPS-loss (M, complete model), the optimization effect of deep NPS-loss on noise texture distribution can be more intuitively observed. Fig. 10(b) presents the radial NPS comparison of the 32×32 ROI area, showing that the complete model with deep NPS loss aligns more closely with the NDCT image in terms of frequency distribution and amplitude. This indicates that the loss effectively promotes noise suppression and balanced texture distribution, demonstrating the positive impact of our deep NPS loss on noise optimization.
Figure 10: (a) Slice of patient L506 from the Mayo dataset shown in Figure 9, with the red dashed box indicating the uniform ROI used for NPS analysis. (b) Radial NPS comparison of this ROI, showing the difference in noise distribution between the model without deep NPS loss (A1) and the complete model with deep NPS loss (M).
IV CONCLUSIONS
In this paper, we propose a hybrid convolutional state-space model, CT-Mamba, for LDCT image denoising. The model combines the multi-scale analysis capability of wavelet transform, the powerful local feature extraction ability of CNN and the long-range dependent modeling advantage of Mamba, so that it can fully capture the details and global information in LDCT images. We propose a new CZSS module, which adopts a spatially coherence “Z” Scan method, which effectively maintains the spatial continuity between adjacent pixels of the image, and further enhances the detail preservation and noise reduction capabilities of the model. Additionally, to the best of our knowledge, this is the first study to use NPS to guide deep learning in LDCT denoising tasks. We designed a deep NPS-loss driven by Mamba, aiming to ensure that the denoised image accurately restores the noise texture distribution of the NDCT image as much as possible, thereby improving the overall image quality and diagnostic value. We evaluated CT-Mamba on two different datasets, and the experimental results show that our model performs excellently in denoising effectiveness, noise texture preservation, and radiomic feature restoration, showing potential to become a representative method of the Mamba framework for LDCT denoising tasks.

### v2

@@
 Abstract
-Low-dose CT (LDCT) significantly reduces the radiation dose received by patients, thereby decreasing potential health risks.
-However, dose reduction introduces additional noise and artifacts, adversely affecting image quality and clinical diagnosis.
+Low-dose CT (LDCT) significantly reduces the radiation dose received by patients, however, dose reduction introduces additional noise and artifacts.
 Currently, denoising methods based on convolutional neural networks (CNNs) face limitations in long-range modeling capabilities, while Transformer-based denoising methods, although capable of powerful long-range modeling, suffer from high computational complexity.
 Furthermore, the denoised images predicted by deep learning-based techniques inevitably exhibit differences in noise distribution compared to normal-dose CT (NDCT) images, which can also impact the final image quality and diagnostic outcomes.
-In recent years, the feasibility of applying deep learning methods to low-dose CT imaging has been demonstrated, leading to significant achievements.
 This paper proposes CT-Mamba, a hybrid convolutional State Space Model for LDCT image denoising.
-The model combines the local feature extraction advantages of CNNs with Mamba’s global modeling capability, enabling it to capture both local details and global context.
-Additionally, a Mamba-driven deep noise power spectrum (NPS) loss function was designed to guide model training, ensuring that the noise texture of the denoised LDCT images closely resembles that of NDCT images, thereby enhancing overall image quality and diagnostic value.
-Experimental results have demonstrated that CT-Mamba performs excellently in reducing noise in LDCT images, enhancing detail preservation, and optimizing noise texture distribution, while demonstrating statistically similar radiomics features to those of NDCT images (p $>$ 0.05).
+The model combines the local feature extraction advantages of CNNs with Mamba’s strength in capturing long-range dependencies, enabling it to capture both local details and global context.
+Additionally, we introduce an innovative spatially coherent ‘Z’-shaped scanning scheme to ensure spatial continuity between adjacent pixels in the image.
+We design a Mamba-driven deep noise power spectrum (NPS) loss function to guide model training, ensuring that the noise texture of the denoised LDCT images closely resembles that of NDCT images, thereby enhancing overall image quality and diagnostic value.
+Experimental results have demonstrated that CT-Mamba performs excellently in reducing noise in LDCT images, enhancing detail preservation, and optimizing noise texture distribution, and exhibits higher statistical similarity with the radiomics features of NDCT images.
 The proposed CT-Mamba demonstrates outstanding performance in LDCT denoising and holds promise as a representative approach for applying the Mamba framework to LDCT denoising tasks.
+Our code will be made available after the paper is officially published: https://github.com/zy2219105/CT-Mamba/.
 Index Terms:

### v3

@@
 The proposed CT-Mamba demonstrates outstanding performance in LDCT denoising and holds promise as a representative approach for applying the Mamba framework to LDCT denoising tasks.
-Our code will be made available after the paper is officially published: https://github.com/zy2219105/CT-Mamba/.
-Index Terms:
+keywords:
 Low-Dose CT , Denoising , State Space Model , Mamba , Noise Power Spectrum , Radiomics
-I Introduction
-Computed tomography (CT) is an essential imaging technique in clinical practice, providing crucial anatomical information that aids physicians in making appropriate medical decisions.
+[…]
+[label1]organization=School of Physics, Beihang University, city=Beijing, country=China
+[…]
+[label2]organization=Emergency Department of the Third Medical Center, Chinese PLA General Hospital,city=Beijing, country=China
+[…]
+[label3]organization=Hangzhou International Innovation Institute, Beihang University,city=Hangzhou, country=China
+[…]
+[label4]organization=Tianmushan Laboratory,city=Hangzhou, country=China
+1 Introduction
+11footnotetext: Linxuan Li and Wenjia Wei are co-first authors.22footnotetext: Github link: https://github.com/zy2219105/CT-Mamba33footnotetext: Contact Email: zy2219105@buaa.edu.cn, weiwenjia@buaa.edu.cn
+Computed tomography (CT) is an essential imaging technique in clinical practice, providing crucial anatomical information that aids physicians in making appropriate medical decisions [49].
 However, frequent CT scans can significantly increase the radiation dose received by patients, particularly in radiation therapy, where multiple scans are often required to achieve precise tumor localization and minimize damage to organs at risk.
@@
 If these issues are not effectively addressed, they will significantly impact the application of LDCT in various clinical scenarios.
-How to obtain high-quality CT images that are comparable to normal-dose CT (NDCT) images and meet the clinical utilities from low-dose scanning protocols has become a long-standing and practically significant problem in the field of CT imaging.
-To address this issue, several LDCT imaging algorithms have been proposed, which can be roughly divided into the following three categories [2, 3]: sinogram-based preprocessing, iterative reconstruction, and image post-processing.
+How to obtain high-quality CT image‘s that are comparable to normal-dose CT (NDCT) images and meet the clinical utilities from low-dose scanning protocols has become a long-standing and practically significant problem in the field of CT imaging.
+To address this issue, several LDCT imaging algorithms have been proposed, which can be roughly divided into the following three categories [48, 23, 51]: sinogram-based preprocessing, iterative reconstruction, and image post-processing.
 Sinogram-based preprocessing methods directly target the raw data acquisition stage and designs specific filters for processing CT projection data obtained through low-dose X-rays.
@@
 In the field of LDCT denoising, several representative CNN-based models have been proposed [32, 24, 11, 21, 31, 50, 20].
-However, CNN frameworks are limited by their lack of long-range modeling capabilities, making it difficult to capture the global contextual information in images [22].
+However, CNN frameworks are limited by their lack of long-range modeling capabilities, making it difficult to capture the global contextual information in images [47, 52].
 Transformers were originally designed for natural language processing, but have now been successfully adapted for image processing, such as Vision Transformer (ViT) [6] for image recognition and SwinTransformer [27] for a general backbone in various vision tasks.
@@
 Therefore, this paper also utilizes NPS to guide deep learning LDCT denoising tasks, ensuring that the predicted images more accurately restore the noise distribution present in NDCT images, thereby improving overall image quality and diagnostic value.
+Figure 1: (a) The overall architecture of the proposed CT-Mamba. (b) The structure of the Multi-Scale Coherence Mamba architecture (MSC-Mamba) in CT-Mamba.
 The contributions of this study are summarized as follows:
@@
 Using radiomics, the proposed CT-Mamba was evaluated by comparing the statistical distribution of radiomics features in different organs between denoised LDCT images and NDCT images, as well as the mean absolute error of pairwise features.
-Figure 1: (a) The overall architecture of the proposed CT-Mamba. (b) The structure of the Multi-Scale Coherence Mamba architecture (MSC-Mamba) in CT-Mamba.
-II Medthod
-II-A Preliminary: State Space Model
+2 Methodology
+2.1 Preliminary: State Space Model
 The state space model (SSM) is a mathematical description of a dynamic system which maps a one-dimensional input function or sequence $x(t)\in{{\mathbb{R}}^{L}}$ to an output $y(t)\in{{\mathbb{R}}^{L}}$ through hidden states $h(t)\in{{\mathbb{R}}^{N}}$.
@@
 where $\overline{\mathbf{K}}\in{{\mathbb{R}}^{L}}$ represents a structured convolutional kernel, $L$ denotes the length of the input sequence $x$, and * represents the convolution operation.
+Figure 2: The structure of High-Frequency Feature Extraction Network (HFEN).
 The recently proposed Mamba model makes further improvements by introducing a selective scanning mechanism.
@@
 As a result, Mamba can model complex temporal dynamics more effectively, as the model can adapt to the evolving characteristics of the input data.
-II-B Framework Overview
+2.2 Framework Overview
 In this paper, we propose CT-Mamba, a hybrid convolutional state space model designed for LDCT image denoising.
@@
 This enables the network to better capture complex structural features within the image, thereby improving overall performance.
-Figure 2: The structure of High-Frequency Feature Extraction Network (HFEN).
 We designed a Multi-Scale Coherence Mamba architecture (MSC-Mamba) to learn the features of each frequency branch obtained by wavelet transform decomposition, as shown in Fig.
 1(b).
-MSC-Mamba captures and fuses frequency features learned from three different scales (32×32, 16×16, and 8×8) by integrating Coherence Z-Scan State Space Block (CZSS).
+MSC-Mamba captures and fuses frequency features learned from three different scales ($H\times W\times C$, $\frac{H}{2}\times\frac{W}{2}\times 4C$, and $\frac{H}{4}\times\frac{W}{4}\times 16C$) by integrating Coherence Z-Scan State Space Block (CZSS).
 At lower scales, MSC-Mamba effectively captures global feature information, while at higher scales, it focuses more on local details.
@@
 Figure 3: The structure of Frequency Domain Attention Module (FDAM).
-II-C Coherence Z-Scan State Space Block
+Figure 4: Illustration of the structure of Z-SSM. The Z-shaped scanning method is adopted to ensure spatial continuity between each adjacent pixel in the image.
+2.3 Coherence Z-Scan State Space Block
 The Coherence Z-Scan State Space Block (CZSS) is a novel feature extraction unit specifically designed for tasks like medical imaging.
@@
 This design ensures that subtle lesions and detailed information within the image are not overlooked, thereby enhancing the denoising effect of LDCT images.
-Figure 4: Illustration of the structure of Z-SSM. The Z-shaped scanning method is adopted to ensure spatial continuity between each adjacent pixel in the image.
 A residual block with two convolutional layers is added to the front of CZSS can locally maintain the frequency-spatial correlation of input features and efficiently transfer local contextual information, thereby enabling more effective utilization and further optimization of these features, as represented by “Res” in Fig.
@@
 CZSS finds a good balance between computational efficiency and feature extraction capability, allowing for stacking more CZSS blocks under similar depth constraints.
-II-D Deep Noise Power Spectrum Loss
+Figure 5: Detailed structure of the dual-branch Deep NPS Loss and its key components. (a) Structure of the Deep NPS Loss. (b) Detailed structure of the u-Feature Net.
+Figure 6: Schematic diagram of the radial transformation process of a 2D NPS signal.
+2.4 Deep Noise Power Spectrum Loss
 Noise texture is highly sensitive to changes in subjective visual quality.
@@
 This approach ensures that the noise texture in the model’s denoised images more closely resembles that of NDCT images, thereby improving overall image quality and diagnostic value.
-Figure 5: Detailed structure of the dual-branch Deep NPS Loss and its key components. (a) Structure of the Deep NPS Loss. (b) Detailed structure of the u-Feature Net.
-Figure 6: Schematic diagram of the radial transformation process of a 2D NPS signal.
-II-E Overall Objective
+2.5 Overall Objective
 The loss function measures the difference between the model’s predicted output (the denoised LDCT image) and the ground truth (NDCT image).
@@
 Since the pre-trained ResNet50 accepts color images as input, we replicated the grayscale CT images across three channels to fit this input requirement.
+This approach has been validated in LDCT tasks [46].
 we extracted feature maps after its four main stages and used L2 loss to measure the similarity between the CT-Mamba output and the ground truth image.
@@
 where ${{\lambda}_{2}}=({{\gamma}_{1}},{{\gamma}_{2}})$.
-III EXPERIMENT DESIGNS AND RESULTS
-III-A Datasets
-In this work, we used a publicly released patient dataset for the 2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge.
+3 EXPERIMENT DESIGNS AND RESULTS
+3.1 Datasets
+In this work, we used a publicly released patient dataset for the 2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge [30].
 This dataset contains paired low-dose and normal-dose abdominal images with a slice thickness of 1 mm for 10 patients.
@@
 We used data from 9 patients for model training, totaling 5,410 image pairs, while data from the remaining patient (L506) was used for testing model performance, totaling 526 pairs.
-Additionally, using the method in [37], we simulated paired low-dose and normal-dose images for 20 patients from the AMOS dataset (referred to as the Simulator dataset) for model training and validation.
+Additionally, using the method in [40], we simulated paired low-dose and normal-dose images for 20 patients from the AMOS dataset (referred to as the Simulator dataset) for model training and validation [18].
 The low-dose images also correspond to a quarter of the normal dose.
 We used data from 19 patients for training, totaling 1,872 image pairs, and data from 1 patient (AMOS-0033) for testing model performance, totaling 106 pairs.
-III-B Implementation details
+3.2 Implementation details
 The proposed CT-Mamba model is implemented in PyTorch and trained in an environment equipped with an NVIDIA RTX 3090 Ti 24G GPU.
@@
 The training process lasts for 250 epochs, with the final model selected based on the optimal loss achieved.
-Figure 7: A set of slice prediction results and difference images obtained from the Mayo dataset L506 patient.
-The enlarged ROI (blood vessel) marked by the red dashed rectangle is located in the lower right corner of the image.
-Figure 8: A set of slice prediction results and difference images obtained from the Mayo dataset L506 patient.
-The enlarged ROI (low-attenuation lesion, tissue edge) marked by the red dashed rectangle is located in the lower right corner of the image.
-III-C Experiment result and design
+3.3 Experiment result and design
 To demonstrate the effectiveness of the proposed CT-Mamba, we selected several advanced representative methods for comparative experiments, including EDCNN [26], RED-CNN [3], Uformer [41], CTformer [38], and VM-UNet [36].
 EDCNN and RED-CNN are CNN-based methods, Uformer and CTformer are Transformer-based methods, and VM-UNet is based on the Mamba framework.
-III-C1 Visual evaluation
+3.3.1 Visual evaluation
 This section presents the visual validation results of the CT-Mamba model compared to various baseline methods.
@@
 To further assess detail representation, an enlarged view of the region of interest (ROI) which is marked by a red dashed rectangle is provided in the lower right corner of each image, along with the PSNR/SSIM quantitative results for each method displayed in the lower left corner.
+Figure 7: A set of slice prediction results and difference images obtained from the Mayo dataset L506 patient.
+The enlarged ROI (blood vessel) marked by the red dashed rectangle is located in the lower right corner of the image.
+Figure 8: A set of slice prediction results and difference images obtained from the Mayo dataset L506 patient.
+The enlarged ROI (low-attenuation lesion, tissue edge) marked by the red dashed rectangle is located in the lower right corner of the image.
 From the results shown in Fig.
@@
 8.
+Figure 9: A set of slice prediction results and difference images obtained from the Simulator dataset AMOS-0033 patient, with the difference images located in the upper right corner.
 In Fig.
@@
 The upper right corner of each predicted image shows the corresponding difference image.
-Figure 9: A set of slice prediction results and difference images obtained from the Simulator dataset AMOS-0033 patient, with the difference images located in the upper right corner.
 We set the window level of the predicted results to 40 HU and the window width to 400 HU (i.e., a range from -160 HU to 240 HU).
 For the difference images, the window level is set to 0 HU with a window width of 200 HU (i.e., a range from -100 HU to 100 HU).
-III-C2 Quantitative analysis
+Table 1: The average quantitative results of different methods for patient(L506) in the Mayo test dataset.
+Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ VIF $\uparrow$
+LDCT 42.3509 0.9471 31.9809 0.2965
+EDCNN 46.6959 0.9819 19.1911 0.3578
+REDCNN 46.1284 0.9799 20.5049 0.3406
+Uformer 44.4342 0.9655 24.8988 0.3192
+CTformer 46.2254 0.9770 20.3881 0.3471
+VM-Unet 46.5672 0.9824 19.4795 0.3606
+CT-Mamba 47.2825 0.9842 17.9850 0.3735
+Table 2: The average quantitative results of different methods for patient(AMOS-0033) in the Simulator test dataset.
+Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ VIF $\uparrow$
+LDCT 38.7673 0.9099 35.5206 0.3573
+EDCNN 50.3082 0.9931 9.2005 0.5186
+REDCNN 49.7609 0.9921 9.8209 0.5098
+Uformer 44.9416 0.9795 17.1556 0.4241
+CTformer 43.1905 0.9673 21.4047 0.4063
+VM-Unet 48.0132 0.9917 11.9833 0.4880
+CT-Mamba 51.2054 0.9942 8.2936 0.5444
+3.3.2 Quantitative analysis
 Our method not only demonstrates significant optimization effects in subjective visual evaluation but also achieves excellent performance across multiple quantitative metrics.
@@
 In Table.
-I and Table.
-II, we present the average quantitative results for PSNR, SSIM, RMSE, and VIF for patient L506 from the Mayo dataset and patient AMOS-0033 from the Simulator dataset, respectively.
+1 and Table.
+2, we present the average quantitative results for PSNR, SSIM, RMSE, and VIF for patient L506 from the Mayo dataset and patient AMOS-0033 from the Simulator dataset, respectively.
 A comprehensive evaluation of the PSNR, SSIM, RMSE, and VIF metrics shows that CT-Mamba model outperforms the comparison methods across all quantitative indicators.
-TABLE I: The average quantitative results of different methods for patient(L506) in the Mayo test dataset.
-Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ VIF $\uparrow$
-LDCT 42.3509 0.9471 31.9809 0.2965
-EDCNN 46.6959 0.9819 19.1911 0.3578
-REDCNN 46.1284 0.9799 20.5049 0.3406
-Uformer 44.4342 0.9655 24.8988 0.3192
-CTformer 46.2254 0.9770 20.3881 0.3471
-VM-Unet 46.5672 0.9824 19.4795 0.3606
-CT-Mamba 47.2825 0.9842 17.9850 0.3735
-TABLE II: The average quantitative results of different methods for patient(AMOS-0033) in the Simulator test dataset.
-Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ VIF $\uparrow$
-LDCT 38.7673 0.9099 35.5206 0.3573
-EDCNN 50.3082 0.9931 9.2005 0.5186
-REDCNN 49.7609 0.9921 9.8209 0.5098
-Uformer 44.9416 0.9795 17.1556 0.4241
-CTformer 43.1905 0.9673 21.4047 0.4063
-VM-Unet 48.0132 0.9917 11.9833 0.4880
-CT-Mamba 51.2054 0.9942 8.2936 0.5444
-TABLE III: The p-values calculated using the Wilcoxon rank-sum test for each model. AP-value of under 0.05 indicates statistical difference. Pairs with statistical differences are marked by *.
-Method Aorta Right Kidney Liver Stomach Small Bowel Urinary Bladder
-LDCT 0.2774 0.0787 0.1626 0.7841 0.7116 0.6320
-EDCNN 0.0557 0.0557 0.3793 0.4701 0.7519 0.5830
-REDCNN 0.0979 0.4080 0.7959 0.9191 0.9893 0.6018
-Uformer 0.0979 0.1208 0.0386* 0.2137 0.4772 0.1766
-CTformer 0.1960 0.1626 0.3544 0.6198 0.6498 0.3040
-VM-Unet 0.0016* 0.3519 0.8361 0.9801 0.8286 0.6741
-CT-Mamba 0.5014 0.7960 0.6872 0.8948 0.8857 0.8032
-TABLE IV: Number of similarity ratio which are higher for each model compared to the LDCT volumes out of the 93 features.
-Method Aorta Right Kidney Liver Stomach Small Bowel Urinary Bladder
-EDCNN 6 1 6 32 61 8
-REDCNN 5 1 5 33 81 12
-Uformer 10 5 5 62 76 10
-CTformer 8 1 8 36 63 7
-VM-Unet 9 5 6 23 67 13
-CT-Mamba 4 2 7 11 4 10
-III-C3 Radiomics reserach
+3.3.3 Radiomics reserach
 Radiomics holds the potential to transform digital medical images into quantitative features that reveal underlying pathology, with promising applications in tumor classification and patient outcome prediction [44].In this study, the TotalSegmentator [42] extension in 3D Slicer with a pretrained nnU-Net [17] was utilized for automatic segmentation of target organs (aorta, right kidney, liver, stomach, small bowel, and urinary bladder), followed by manual adjustments to ensure accuracy.
@@
 For each volume, the Wilcoxon rank-sum test was employed to evaluate the consistency of radiomics features distribution between the denoised images and the reference NDCT (p $>$ 0.05).
-In Table III, we show the p-values for each model across the target organs (aorta, right kidney, liver, stomach, small bowel, and urinary bladder).
+In Table 3, we show the p-values for each model across the target organs (aorta, right kidney, liver, stomach, small bowel, and urinary bladder).
 A higher p-value indicates greater similarity to the radiomics features distribution of the NDCT.
-As shown in Table III, while VM-UNet performed well for the liver and stomach, its performance for the aorta was poor, exhibiting significant differences in radiomics features distributios compared to the NDCT.
+As shown in Table 3, while VM-UNet performed well for the liver and stomach, its performance for the aorta was poor, exhibiting significant differences in radiomics features distributios compared to the NDCT.
 In contrast, the proposed CT-Mamba model demonstrated superior performance across multiple organs, including the aorta, right kidney, and urinary bladder.
@@
 For each model, the $R$ across the target organs were compared against $R_{\text{LDCT}}$, and the number of features with $R$ exceeding $R_{\text{LDCT}}$ was recorded.
-As shown in Table IV, the proposed CT-Mamba demonstrated superior performance across all target organs.
+As shown in Table 4, the proposed CT-Mamba demonstrated superior performance across all target organs.
 Notably, for the small bowel, where other models performed poorly, CT-Mamba achieved the best performance, further highlighting its robustness and effectiveness.
 The radiomics analysis indicate that the proposed CT-Mamba effectively preserves the shape and texture features of multiple organs within the LDCT images, demonstrating its potential value for clinical applications.
-III-C4 Ablation study
+Table 3: The p-values calculated using the Wilcoxon rank-sum test for each model. AP-value of under 0.05 indicates statistical difference. Pairs with statistical differences are marked by *.
+Method Aorta Right Kidney Liver Stomach Small Bowel Urinary Bladder
+LDCT 0.2774 0.0787 0.1626 0.7841 0.7116 0.6320
+EDCNN 0.0557 0.0557 0.3793 0.4701 0.7519 0.5830
+REDCNN 0.0979 0.4080 0.7959 0.9191 0.9893 0.6018
+Uformer 0.0979 0.1208 0.0386* 0.2137 0.4772 0.1766
+CTformer 0.1960 0.1626 0.3544 0.6198 0.6498 0.3040
+VM-Unet 0.0016* 0.3519 0.8361 0.9801 0.8286 0.6741
+CT-Mamba 0.5014 0.7960 0.6872 0.8948 0.8857 0.8032
+Table 4: Number of similarity ratio which are higher for each model compared to the LDCT volumes out of the 93 features.
+Method Aorta Right Kidney Liver Stomach Small Bowel Urinary Bladder
+EDCNN 6 1 6 32 61 8
+REDCNN 5 1 5 33 81 12
+Uformer 10 5 5 62 76 10
+CTformer 8 1 8 36 63 7
+VM-Unet 9 5 6 23 67 13
+CT-Mamba 4 2 7 11 4 10
+3.3.4 Ablation study
 To evaluate the effectiveness of the key components of CT-Mamba, we designed three ablation experiments and conducted quantitative validation on patient L506 from the Mayo dataset:
 (A1) Effectiveness of deep NPS loss: Excluding deep NPS-loss during training.
-(A2) Effectiveness of the Frequency Domain Attention Module: Excluding the FDAM module.
 (A2) Effectiveness of the Scanning Method: Replacing the Coherence ”Z” Scan with the classic Visual Mamba scanning method.
@@
 The quantitative results of each ablation experiment are summarized in Table.
-V.
+5.
 The results indicate that the complete CT-Mamba model achieves the best performance across all metrics, further validating the significant contribution of each component to the overall performance of model.
-TABLE V: Quantitative comparison of different ablation experiments for patient(L506) in the Mayo test dataset.
+Table 5: Quantitative comparison of different ablation experiments for patient(L506) in the Mayo test dataset.
 Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ VIF $\uparrow$
@@
 A1 47.0101 0.9830 18.5470 0.3679
-A2 47.2301 0.9841 18.0940 0.3723
-A3 47.2245 0.9841 18.1040 0.3728
+A2 47.2245 0.9841 18.1040 0.3728
 M 47.2825 0.9842 17.9850 0.3735
@@
 This indicates that the loss effectively promotes noise suppression and balanced texture distribution, demonstrating the positive impact of our deep NPS loss on noise optimization.
-Figure 10: (a) Slice of patient L506 from the Mayo dataset shown in Figure 9, with the red dashed box indicating the uniform ROI used for NPS analysis.
+Figure 10: (a) Slice of patient L506 from the Mayo dataset shown in Fig.
+8, with the red dashed box indicating the uniform ROI used for NPS analysis.
 (b) Radial NPS comparison of this ROI, showing the difference in noise distribution between the model without deep NPS loss (A1) and the complete model with deep NPS loss (M).
-IV CONCLUSIONS
+4 CONCLUSIONS
 In this paper, we propose a hybrid convolutional state-space model, CT-Mamba, for LDCT image denoising.
@@
 We evaluated CT-Mamba on two different datasets, and the experimental results show that our model performs excellently in denoising effectiveness, noise texture preservation, and radiomic feature restoration, showing potential to become a representative method of the Mamba framework for LDCT denoising tasks.
+CRediT authorship contribution statement
+Linxuan Li: Writing - Original Draft, Methodology, Software, Conceptualization, Investigation.
+Wenjia Wei: Writing - Original Draft, Software, Methodology.
+Luyao Yang: Writing - Original Draft, Supervision, Validation.
+Wenwen Zhang: Visualization.
+Jiashu Dong: Data curation.
+Yahua Liu: Writing – review and editing.
+Wei Zhao: Writing – review and editing, Funding acquisition, Methodology, Conceptualization, Investigation.
+Acknowledgments
+This work was supported in part by the Natural Science Foundation of Zhejiang Province (No.
+LZ23A050002), Natural Science Foundation of Beijing (No.
+L246073), Natural Science Foundation of China (No.
+12175012), the “111 center” (No.
+B20065), and the Fundamental Research Funds for the Central Universities, China.